You shoot an arrow straight up at 50 m/s. When will it run out of speed?
Answer #1
You shoot an arrow straight up at 50 m/s. When will it run out of speed? ------------ h(t) = -16y^2+50t It will "run out of steam" when it's height is maximum. That occurs when t = -b/2a = -50/[2(-16)] = 50/32 = 1.563 seconds ========================== Cheers, Stan H.
answered by: layth